# IEEEexample:BSTcontrol

### v1

@@ 3.4 Dataset Statistics
The DrRank dataset, constructed through the aforementioned process, includes 30 disease types, comprising 23 cancers and 7 other critical illnesses. Each disease is linked to one or more treatment options such as medication, surgery, radiotherapy, and, in the case of leukemia, Hematopoietic Stem Cell Transplantation (HSCT). Although most diseases are associated with a single treatment option, some involve multiple treatments, resulting in a total of 38 unique disease-treatment pairs. Our dataset includes 4325 doctor profiles, of which 2,593 are unique. Some doctors are listed in multiple disease-treatment pairs due to their expertise spanning a variety of conditions and treatments. Fig. 4 illustrates the unbalanced distribution of annotated professional relevance scores among positive samples across all the disease-treatment pairs in the DrRank dataset. Furthermore, Fig. 3 reveals substantial variability in the length of doctor profiles, with the average length in the DrRank dataset exceeding that of our initial doctor pool. These findings underscore the complexity of the doctor ranking task in real-world scenarios, given the significant variation in the volume and diversity of available information. Looking ahead, we plan to expand the dataset by incorporating more disease-treatment pairs.
4 LLM-based Doctor Ranking Framework
@@ 4.1 Problem Formulation
Definition 1 (Doctor Recommendation)
Given a patient’s medical need $q\in\mathcal{Q}$ and a candidate doctor $d\in\mathcal{D}$, the goal of doctor recommendation is to evaluate the professional relevance of doctors to patients’ specific medical needs, i.e., $f(q,d)\rightarrow\mathbb{R}$, which serves as the basis for doctor ranking.
@@ 4.2 LLM-based Doctor Ranking
Fig. 1 presents a comprehensive overview of our proposed doctor ranking framework, which integrates several key components to ensure accurate, efficient, and explainable doctor recommendation. Below, we will elaborate on the details of our framework.
@@ 4.2.1 Exploiting Efficient and Effective Pointwise Ranking
Although pairwise and listwise ranking methods have the capability to compare candidates relatively during inference time, their practical application in real-world scenarios remains extremely challenging. Pairwise comparisons suffer from quadratic complexity [13], incurring significant computational overhead. Meanwhile, the sliding window mechanism employed by listwise method is inherently difficult to parallelize [12], resulting in substantial latency in large-scale medical data sorting tasks. Additionally, both pairwise and listwise methods are susceptible to position bias [26, 27, 13], where the order of candidates can influence the ranking outcomes. This bias introduces instability, severely compromising the credibility of ranking results, particularly in critical medical scenarios where precision and reliability are paramount. In contrast, the pointwise ranking method excels with its high degree of parallelism, allowing for independent scoring of each candidate [23, 24]. A notable advantage of the pointwise method lies in the fact that it generates scores which are free from position bias. This ensures that the scores are stable, facilitating easy reproducibility of ranking results. Furthermore, by harnessing the robust language generation capabilities of LLMs, generating the scoring rationale for each candidate becomes straightforward, enhancing the transparency and trustworthiness of the ranking process. In comparison, pairwise and listwise methods usually focus on relative comparisons between pairs or lists of candidates, making it difficult to extract and present individual, easily understandable rationales for each candidate. Considering the aspects of parallel efficiency, score stability and interpretability mentioned above, the pointwise ranking method emerges as the ideal choice for doctor ranking.
@@ 4.2.2 Utilizing Fine-grained Labels to Optimize Distinction Relevance
Existing research utilizing LLMs as zero-shot pointwise rankers can be broadly categorized into two approaches: relevance generation [24, 11] and query generation [23]. This work focuses on the relevance generation approach. In this approach, the LLM is prompted to determine whether a document is relevant to a query, using the logits of “Yes” and “No” tokens to derive a relevance score. However, studies involving human subjects have demonstrated that binary options can sometimes lead to biased responses, whereas offering options with a higher level of granularity can produce more dependable outcomes [28]. Zhuang et al. [11] suggest incorporating detailed relevance labels into the prompt for pointwise LLM rankers, allowing them to better distinguish among documents with differing degrees of relevance to the query and thus generate a more precise ranking. Inspired by this, we employ a set of 5-level graded labels [“Top”, “High”, “Mid”, “Low”, “Not Relevant”] to assess the professional relevance of doctors in addressing patients’ specific medical needs related to a particular disease and its treatment options. For each query-doctor pair $(q,{d_{i}})$, we instruct the LLM to evaluate their professional relevance by selecting from the provided graded labels. We can obtain the corresponding $logit_{k}$ for each graded label $l_{k}$ from the LLM output as (1), which represents the raw confidence values of evaluation.
$logit_{i,k}=\mathrm{LLM}(l_{k}|q,d_{i},prompt),\;k\in[0,1,2,3,4]$ (1)
@@
We can use these logits to derive the ranking score by calculating the expected value. Firstly, we need to assign a score $s_{k}$ to each graded label $l_{k}$ Through empirical investigation, we find that naively assigning $s_{k}=k$ (with 0 to 4 ordered from least to most relevance) already yields excellent performance. Then, we can calculate the expected scoring value as follows:
$f(q,d_{i})=\sum_{k}prob_{i,k}\cdot s_{k},$ (2)
@@
where $\text{where }prob_{i,k}=\frac{\exp(logit_{i,k})}{\sum_{k^{\prime}}\exp(logit_{i,k^{\prime}})}.$
In this context, the marginal probability $prob_{i,k}$ can be interpreted as the likelihood of the doctor $d_{i}$’s relevance belonging to the graded label $l_{k}$. In fact, this approach to finding the expected average is akin to sampling the outputs of LLMs multiple times [11]. In preliminary experiments, we surprisingly observed that the first token output by instruction-tuned LLMs does not always match our predefined 5-level graded labels, complicating the extraction of logits for deriving ranking scores. Therefore, we propose a forced elicitation mechanism, which involves appending a predefined labeling prefix to elicit the desired graded labels from the LLM. The labeling prefix (i.e., “The professional relevance of the candidate doctor is ”) ensures that the first token output by the LLM matches our expectations perfectly.
@@ 4.2.3 Generating Ranking Criteria as Standard Guidelines
LLMs are widely known as extremely intricate “black-box” systems [29]. The iterative nature of next-token prediction results in ambiguous assessment criteria for zero-shot pointwise rankers [30], leading to a lack of consistency and transparency in the scoring process. Establishing clear criteria by which LLMs score doctors is crucial for building an accurate and explainable doctor recommendation system. Recent studies on human annotation practices have emphasized the importance of a standardized annotation guideline for ensuring clear and consistent communication with annotators [31, 32]. Inspired by this, we argue that standard ranking criteria should be provided for LLMs to ensure doctors are evaluated under a unified and transparent standard, thereby enhancing the consistency and accuracy of the final doctor ranking results. To this end, we meticulously design a prompt (as shown in Fig. 5) to instruct LLMs to generate relevant doctor ranking criteria for specific disease-treatment pairs. Initially, we repeatedly prompt LLMs to generate multiple sets of ranking criteria for a specific disease-treatment pair. Subsequently, we manually inspect these criteria sets to select the best one, based on factors such as comprehensiveness, practicality, and overall format. Using the optimal criteria as an example, we adopt a one-shot prompting approach to generate consistent ranking criteria for other disease-treatment pairs. Our method dynamically integrates query-specific (i.e., disease and treatment) ranking criteria into the ranking prompt, serving as contextual guidelines for the rankers’ decisions. These query-specific factors are essential for aligning ranking results with patients’ needs and ensuring that ranking decisions are based on the most relevant professional factors. Recently some researchers use multi-perspective criteria to further enhance pointwsie rankers [30]. This approach can be viewed as a more complex variant of our method, involving multiple rounds of LLM inference and result ensembling from each perspective. While it potentially offers enhanced performance, it results in increased complexity and computational cost. Our method, in contrast, offers a simpler yet effective solution. We verify the consistent positive impact of generated criteria on the doctor ranking performance of LLMs in our ablation experiments.
Figure 5: Prompt for ranking criteria generation.
4.2.4 Generating Evaluation Rationales as Recommendation Explanations
@@ 5 Experiments
In this section, we evaluate the proposed framework on the DrRank dataset under a zero-shot setting. Meanwhile, we present various discussions on doctor recommendation.
@@ 5.1 Experimental Settings
We omit the dataset description in this section since it has been introduced in Section 3. Other experimental settings will be described in the following parts.
Table 1: The performance and efficiency of each model on the DrRank dataset. The model sizes and maximum context lengths for cross-encoder models are indicated as [parameters, tokens]. The maximum context and generation length of Qwen2.5 series are 32K and 8K respectively. “BS” is an abbreviation for “batch size”. Scoring Time and Explanation Time are the average time consumption of each batch. Average results that are significantly (paired t-test, $p<0.05$ ) better than without ranking criteria are marked with *. The top three performances of the models with ranking criteria are highlighted in bold. The top three performances of the models without ranking criteria are underlined.
Model Running Info NDCG@10 PNR Scoring Time (s) Explanation Time (s)
Cross-Encoder
jinaai/jina-reranker-v2-base-multilingual [278M, 1K] 8 BS, 1 GPU 56.91 1.79 0.01 -
Alibaba-NLP/gte-multilingual-reranker-base [306M, 8K] 8 BS, 1 GPU 71.96 4.01 0.01 -
BAAI/bge-reranker-v2-m3 [568M, 8K] 8 BS, 1 GPU 69.26 3.42 0.02 -
LLM-based Pointwise Ranker
General-purpose LLM
Qwen2.5-0.5B-Instruct w/ Ranking Criteria 5 BS, 1 GPU 49.85 1.68 0.02 6.36
Qwen2.5-0.5B-Instruct w/o Ranking Criteria 10 BS, 1 GPU 56.21 1.94 0.02 5.22
Qwen2.5-1.5B-Instruct w/ Ranking Criteria 15 BS, 1 GPU 62.33 3.38 0.09 13.82
Qwen2.5-1.5B-Instruct w/o Ranking Criteria 15 BS, 1 GPU 62.04 3.28 0.11 9.49
Qwen2.5-3B-Instruct w/ Ranking Criteria 4 BS, 1 GPU 67.48 4.23 0.11 13.4
Qwen2.5-3B-Instruct w/o Ranking Criteria 5 BS, 1 GPU 66.95 4.12 0.11 8.55
Qwen2.5-7B-Instruct w/ Ranking Criteria 6 BS, 1 GPU 74.35* 5.31* 0.23 22.62
Qwen2.5-7B-Instruct w/o Ranking Criteria 8 BS, 1 GPU 72.68 4.9 0.16 10.51
Qwen2.5-7B-Instruct w/ Random Ranking Criteria 8 BS, 1 GPU 72.52 4.6 0.23 23.22
Qwen2.5-14B-Instruct w/ Ranking Criteria 8 BS, 2 GPUs 75.63* 6.87* 1.48 32.63
Qwen2.5-14B-Instruct w/o Ranking Criteria 12 BS, 2 GPUs 73.56 6.21 1.73 28.68
Qwen2.5-14B-Instruct w/ Random Ranking Criteria 8 BS, 2 GPUs 75.51 5.27 1.46 28.8
Qwen2.5-32B-Instruct w/ Ranking Criteria 8 BS, 4 GPUs 77.41 6.13 0.1 60.21
Qwen2.5-32B-Instruct w/o Ranking Criteria 8 BS, 4 GPUs 77.19 6.1 0.1 36.42
Qwen2.5-32B-Instruct w/ Random Ranking Criteria 8 BS, 4 GPUs 76.31 5.44 0.1 57.49
Qwen2.5-72B-Instruct w/ Ranking Criteria 5 BS, 3 GPUs 75.93 5.38 4.84 110.08
Qwen2.5-72B-Instruct w/o Ranking Criteria 10 BS, 4 GPUs 75.41 5.38 5.05 82.26
Medical LLM
MING-MOE-7B w/ Ranking Criteria 6 BS, 1 GPU 48.79 1.19 1.63 41.25
MING-MOE-7B w/o Ranking Criteria 10 BS, 1 GPU 45.43 1.25 2.14 34.22
@@ 5.1.1 Baseline Methods
We selected three popular open-sourced cross-encoder models as baselines, namely BAAI/bge-reranker-v2-m3555https://huggingface.co/BAAI/bge-reranker-v2-m3 [36], Alibaba-NLP/gte-multilingual-reranker-base, jinaai/jina-reranker-v2-base-multilingual [^6] https://huggingface.co/jinaai/jina-reranker-v2-base-multilingual, which have demonstrated state-of-the-art effectiveness in various out-of-domain tasks. We concatenate pairs of the patient query and detailed text constructed from the doctor profile, and then pass these concatenated pairs through the models to directly generate a relevance score. Importantly, these cross-encoder models are typically trained on general domain text ranking datasets, such as MS MARCO [37]. Since they have likely not encountered doctor ranking tasks during their training phase, we believe that the experimental results will rely entirely on the generalization ability of these models in the context of doctor ranking.

### v2

@@ 3.4 Dataset Statistics
The DrRank dataset, constructed through the aforementioned process, includes 30 disease types, comprising 23 cancers and 7 other critical illnesses. Each disease is linked to one or more treatment options such as medication, surgery, radiotherapy, and, in the case of leukemia, Hematopoietic Stem Cell Transplantation (HSCT). Although most diseases are associated with a single treatment option, some involve multiple treatments, resulting in a total of 38 unique disease-treatment pairs. Our dataset includes 4325 doctor profiles, of which 2,593 are unique. Some doctors are listed in multiple disease-treatment pairs due to their expertise spanning a variety of conditions and treatments. Fig. 3 illustrates the distribution of annotated professional relevance scores among positive samples across all the disease-treatment pairs in the DrRank dataset, where the x-axis represents the professional relevance (ranging from 1 to 5) and the y-axis lists various disease-treatment pairs. Furthermore, in Fig. 4, the x-axis represents the length of tokenized doctor profiles, and the y-axis represents the cumulative percentage of sequences. The vertical dashed lines at 1024, 2048, and 4096 tokens indicate key truncation thresholds. Each of these thresholds is annotated with the percentage of the DrRank dataset that is covered up to that token count. The distribution reveals substantial variability in the length of doctor profiles, with the average length in the DrRank dataset exceeding that of our initial doctor pool. These findings underscore the complexity of the doctor ranking task in real-world scenarios, given the significant variation in the volume and diversity of available information.
Fig.  4: Cumulative distribution functions (CDF) of doctor profile lengths from two datasets: DrRank and Collected Doctor Pool.
4 Methodology
@@ 4.1 Problem Formulation
Definition 1 (Doctor Ranking)
Given a patient’s medical need $q\in\mathcal{Q}$ and a candidate doctor $d\in\mathcal{D}$, the goal of doctor ranking is to evaluate the professional relevance of doctors to patients’ specific medical needs, i.e., $f(q,d)\rightarrow\mathbb{R}$.
@@ 4.2 LLM-based Doctor Ranking
Fig. 1 presents a comprehensive overview of our proposed ranking framework, which integrates several key components to ensure accurate, efficient, and explainable doctor ranking. Below, we will elaborate on the details of our framework.
@@ 4.2.1 Exploiting Efficient and Effective Pointwise Ranking
Although pairwise and listwise ranking methods have the capability to compare candidates relatively during inference time, their practical application in real-world scenarios remains extremely challenging. Pairwise comparisons suffer from quadratic complexity [13], incurring significant computational overhead. Meanwhile, the sliding window mechanism employed by listwise method is inherently difficult to parallelize [12], resulting in substantial latency in large-scale data sorting tasks. Additionally, both pairwise and listwise methods are susceptible to position bias [30, 31, 13], where the order of candidates can influence the ranking outcomes. This bias introduces instability, severely compromising the credibility of ranking results, particularly in critical medical scenarios where precision and reliability are paramount.
In contrast, the pointwise ranking method excels with its high degree of parallelism, allowing for independent scoring of each candidate [22, 23]. A notable advantage of the pointwise method lies in the fact that it generates scores which are free from position bias. This ensures that the scores are stable, facilitating easy reproducibility of ranking results. Furthermore, by harnessing the robust language generation capabilities of LLMs, generating the scoring rationale for each candidate becomes straightforward, enhancing the transparency and trustworthiness of the ranking process. In comparison, pairwise and listwise methods usually focus on relative comparisons between pairs or lists of candidates, making it difficult to extract and present individual, easily understandable rationales for each candidate. Considering the aspects of parallel efficiency, score stability and interpretability, the pointwise ranking method emerges as the ideal choice for doctor ranking.
@@ 4.2.2 Utilizing Fine-grained Labels to Optimize Distinction Relevance
Existing research utilizing LLMs as zero-shot pointwise rankers can be broadly categorized into two approaches: query generation [22] and relevance generation [23, 11]. This work focuses on the relevance generation approach. In this approach, the LLM is prompted to determine whether a document is relevant to a query, using the logits of “Yes” and “No” tokens to derive a relevance score. However, studies involving human subjects have demonstrated that binary options can sometimes lead to biased responses, whereas offering options with a higher level of granularity can produce more dependable outcomes [32]. Zhuang et al. [11] suggest incorporating detailed relevance labels into the prompt for pointwise LLM rankers, allowing them to better distinguish among documents with differing degrees of relevance to the query and thus generate a more precise ranking.
Inspired by these insights and aligned with the data‐annotation labels (Section 3.2), we adopt a five‐level grading scheme—“Top”, “High”, “Mid”, “Low”, and “Not Relevant”—to evaluate each doctor’s professional relevance for addressing patients’ specific medical needs. For each query-doctor pair $(q,{d_{i}})$, we instruct the LLM to evaluate their professional relevance by selecting from the provided graded labels, with the prompt $p$ shown in Fig. 1. We can obtain the corresponding $\mathrm{logit_{k}}$ for each graded label $l_{k}$ from the LLM output as (1), which represents the raw confidence values of evaluation.
$\begin{split}\mathrm{logit_{i,k}}&=\mathrm{LLM}(l_{k}\mid p,q,d_{i}),\\ \text{where }k&\in\{0,1,2,3,4\}.\end{split}$ (1)
@@
We can use these logits to derive the ranking score by calculating the expected value. Firstly, we need to assign a score $s_{k}$ to each graded label $l_{k}$ . Through empirical investigation, we find that naively assigning $s_{k}=k$ (with 0 to 4 ordered from least to most relevance) already yields excellent performance. Then, we can calculate the expected scoring value as follows:
@@
$\begin{split}f(q,d_{i})&=\sum_{k}\mathrm{prob_{i,k}}\cdot s_{k},\\ \text{where }\mathrm{prob_{i,k}}&=\frac{\exp(\mathrm{logit_{i,k}})}{\sum_{k^{\prime}}\exp(\mathrm{logit_{i,k^{\prime}}})}.\end{split}$ (2)
In this context, the marginal probability $\mathrm{prob_{i,k}}$ can be interpreted as the likelihood of the doctor $d_{i}$’s relevance belonging to the graded label $l_{k}$. In fact, this approach to finding the expected average is akin to sampling the outputs of LLMs multiple times [11].
In preliminary experiments, we observed an unexpected behavior: instruction-tuned LLMs, when prompted to generate a graded assessment, do not always produce outputs that directly correspond to our predefined 5-level label set. Specifically, the first token or phrase generated by the model may deviate from the expected label space, introducing ambiguity and inconsistency when extracting the associated logits for subsequent ranking score computations. This inconsistency undermines the reliability of pointwise scoring mechanisms that rely on matching model outputs to specific labels. To address this issue, we propose a forced elicitation mechanism. This mechanism involves appending a fixed, semantically aligned labeling prefix to the end of each prompt to steer the model’s generation toward a constrained, predictable output space. For instance, by appending the phrase “The professional relevance of the candidate doctor is ”, we provide a strong inductive bias that encourages the model to follow with one of the target graded labels. Empirically, we find that this prefix effectively anchors the model’s generation, ensuring that the first generated token aligns exactly with our predefined label set. This alignment not only facilitates accurate logit extraction but also enhances the interpretability and consistency of model behavior across queries. This forced elicitation strategy can be seen as a lightweight yet robust form of constrained decoding, implemented at the prompt design level without requiring modification of the LLM’s architecture or decoding algorithm.
@@ 4.2.3 Generating Ranking Criteria as Standard Guidelines
LLMs are widely known as extremely intricate “black-box” systems [33]. The iterative nature of next-token prediction results in ambiguous assessment criteria for zero-shot pointwise rankers [28], leading to a lack of consistency and transparency in the scoring process. Establishing clear criteria by which LLMs score doctors is crucial for building an accurate and explainable doctor ranking system. Recent studies on human annotation practices have emphasized the importance of a standardized annotation guideline for ensuring clear and consistent communication with annotators [34, 35].
Inspired by these, we argue that standard ranking criteria should be provided for LLMs to ensure doctors are evaluated under a unified and transparent standard, thereby enhancing the consistency and accuracy of the final doctor ranking results. To this end, we meticulously design a prompt (as shown in Fig. 5) to instruct LLMs to generate doctor ranking criteria for specific disease-treatment pairs. Initially, we repeatedly prompt LLMs to generate multiple sets of ranking criteria for a specific disease-treatment pair. Subsequently, we manually inspect these criteria sets to select the best one, based on factors such as comprehensiveness, practicality, and overall format. Using the optimal criteria as an example, we adopt a one-shot prompting approach to generate consistent ranking criteria for other disease-treatment pairs. Our method dynamically integrates disease-specific ranking criteria into the ranking prompt, serving as contextual guidelines for the rankers’ decisions. These disease-specific factors are essential for aligning ranking results with patients’ needs and ensuring that ranking decisions are based on the most relevant professional factors.
Fig.  5: Prompt for ranking criteria generation.
4.2.4 Generating Rationales as Ranking Explanations
In the context of medical applications, particularly doctor ranking, providing transparent and comprehensible explanations is essential. Patients often seek not only a ranked list of candidate doctors, but also an understanding of why a particular doctor is ranked higher than others. Such explanations help alleviate uncertainty and foster trust in the recommendation system [10].
@@ 5.1 Experimental Settings
We evaluate the proposed framework on the DrRank dataset. We omit the dataset description in this section since it has been introduced in Section 3. Other experimental settings will be described in the following parts.
@@ 5.1.1 Baseline Methods
As discussed in Section 2.1, related work in doctor recommendation faces significant evaluation hurdles due to the lack of standard benchmarks. This issue is further compounded by the unavailability of open-source data and models in many existing approaches [6, 8, 9], alongside the intricate and diverse nature of real-world medical scenarios. To advance research in doctor ranking, we contribute a professionally annotated dataset and benchmark a set of general-purpose ranking models, including bi-encoder and cross-encoder architectures, as zero-shot baselines. In the bi-encoder setup, patient queries and doctor profile texts are independently encoded into dense vectors, and their cosine similarity is used as the ranking score. In contrast, the cross-encoder setup concatenates the patient query with structured doctor profile information and feeds the combined sequence into a reranker to compute a relevance score. For the bi-encoder baselines, we include bge-m3 [36], jina-embeddings-v3 [37], e5-mistral-7b-instruct [38], and gte-Qwen2-7B-instruct [39]. For cross-encoders, we adopt jina-reranker-v2-base, bce-reranker-base, bge-reranker-v2-m3, and bge-reranker-v2-gemma. These models have demonstrated strong performance on a variety of out-of-domain retrieval and ranking tasks. Since they are typically trained on general-domain datasets such as MS MARCO [40], their performance on DrRank reflects their zero-shot generalization ability in the healthcare domain. We also report results for pairwise and listwise LLM-based ranking (described in Section 2.2), where each case is evaluated by jointly considering two or more candidate doctors. While pairwise and listwise approaches can deliver improved performance, they come at the expense of greater computational cost and markedly higher inference latency. As such, these results are not directly comparable to pointwise methods. Instead, we present them as an upper bound to illustrate the potential of pairwise and listwise ranking, particularly when leveraging state-of-the-art LLMs such as gpt-4o.
Table 1: Model Performance on the DrRank dataset. Model sizes of Bi-Encoder and Cross-Encoder are denoted as [parameters]. Model with specifically generated ranking criteria is marked with *. Model with randomly assigned ranking criteria is marked with +. The top results in each group are highlighted in bold.
Category Model NDCG@10 Recall@10
Bi-Encoder bge-m3 [568M] 63.49 29.00
jina-embeddings-v3 [572M] 60.54 28.10
e5-mistral-7b-instruct [7.11B] 69.21 29.98
gte-Qwen2-7B-instruct [7.61B] 70.96 30.03
Cross-Encoder jina-reranker-v2-base [278M] 57.98 25.02
bce-reranker-base_v1 [279M] 62.86 28.37
bge-reranker-v2-m3 [568M] 69.36 29.01
bge-reranker-v2-gemma [2.51B] 69.59 29.91
LLM-based Ranker Pairwise Qwen2.5-32B-Instruct 78.13 34.17
Listwise gemini-2.0-flash 81.59 31.81
Qwen2.5-32B-Instruct 82.39 31.86
claude-3-5-sonnet-20241022 83.20 32.07
gpt-4o-2024-08-06 84.24 32.21
Pointwise (Ours) Qwen2.5-0.5B-Instruct* 45.01 22.16
Qwen2.5-0.5B-Instruct 55.18 24.47
Qwen2.5-1.5B-Instruct* 58.49 27.01
Qwen2.5-1.5B-Instruct 59.65 26.91
Qwen2.5-3B-Instruct* 65.22 27.95
Qwen2.5-3B-Instruct 64.28 27.74
Qwen2.5-7B-Instruct* 73.73 29.99
Qwen2.5-7B-Instruct+ 72.16 29.37
Qwen2.5-7B-Instruct 71.44 29.27
Qwen2.5-32B-Instruct* 77.41 30.73
Qwen2.5-32B-Instruct+ 76.61 30.40
Qwen2.5-32B-Instruct 76.57 30.32
Baichuan-M1-14B-Instruct* 74.67 29.82
Baichuan-M1-14B-Instruct 71.93 28.49